Morphine (C17H19NO3), a narcotic used in painkillers, is a weak organic base.

A) If the pH of a 7.0×10−4M solution of morphine is 9.5, what is the value of Kb?

B) What is the value of pKb?

Express your answer using one decimal place.
Answer #1

pH=9.5 polt =14- pH = 14-9.5=4.5 pott - - Log 80 H3 = 45 [OH-] = 10 pot = 1045 =0.0000316 = x initial concentration of morphi